OpenStudy (anonymous):

A sum of money lent at 2.5% per annum yielded a simple interest of Rs. 5000. If Rs. 45, 000 were paid to clear the loan then find the time for which the money was borrowed.

OpenStudy (anonymous):

Please provide steps if possible.

OpenStudy (anonymous):

\[Interest = Principal \times Rate \times Time\]

OpenStudy (anonymous):

\[Amount = Principal \times (1 + Rate \times Time)\]

OpenStudy (anonymous):

Thanks, I never knew that formula